OpenStudy (anonymous):

Solve the equation and check for extraneous solutions. Click to see question!

OpenStudy (anonymous):

\[\sqrt{7x+2}-2=7x\]

OpenStudy (cwrw238):

\[\sqrt{7x+2} = 7x + 2\]

OpenStudy (cwrw238):

so what must 7x+ 2 be?

OpenStudy (anonymous):

Huuhh?

OpenStudy (anonymous):

Is that the answer?

OpenStudy (cwrw238):

what number is square root of itself?

OpenStudy (cwrw238):

1 right?

OpenStudy (anonymous):

I'm so bad at math and I have no idea what your talking about.

OpenStudy (cwrw238):

7x + 2 = 1 7x = -1 x = -1/7

OpenStudy (anonymous):

Your sure that's right?

OpenStudy (cwrw238):

yea = try plugging in x = -1/7 in the original equation: left hand side of equation: sqrt ( 7 * -1/7 + 2) - 2 = sqrt(-1+2) - 2 = sqrt1 - 2 = -1 right hand side: 7*-1/7 + 2 = 1 - no thats not right

OpenStudy (anonymous):

So what's the answer? haha.

OpenStudy (anonymous):

Add 2 to both sides, square both sides, solve for x. Check your solutions in the original formula. One of them probably won't work

OpenStudy (anonymous):

7x+2=(7x+2)^2 7x+2=49x^2+28x+4 49x^2+21x+2=0

OpenStudy (anonymous):

Use quadratic formula, you'll get 2 answers. Check them both in the original equation and see which one works.

OpenStudy (cwrw238):

my concentration is poor today

OpenStudy (anonymous):

49x^2+21x+2=0 SO this is the answer?

OpenStudy (cwrw238):

oh hold - -1/7 is a solution - i made a mistake when i checked it rhs = -1

OpenStudy (anonymous):

That's not the answer. Use the quadratic formula to get the answers from there.

OpenStudy (anonymous):

Can you please just give me the answer I'm so bad at this and this is my last question!

OpenStudy (cwrw238):

yes -1/7 is correct

OpenStudy (cwrw238):

check back - i made a mistake right hand side = 7 * -1/7 = -1

OpenStudy (cwrw238):

yes it works in original equation

OpenStudy (anonymous):

Which one is the answer?!

OpenStudy (cwrw238):

-1/7 is correct

OpenStudy (anonymous):

Yes, i double checked, there are 2 answers here...

OpenStudy (anonymous):

-1/7 and -2/7 work, which you'd get from the quadratic formula